# A Multi-Fidelity Control Variate Approach for Policy Gradient Estimation

### v3

@@ Appendix D Implementation Details and Hyperparameter Setup
This section provides implementation details and the hyperparameter setup for all evaluated approaches. Details of the off-dynamics task settings described in Section˜6.1 are available in the ODRL benchmark (Lyu et al., 2024b). Table˜1 summarizes the hyperparameter configurations for all evaluated methods. To ensure reproducibility, we release the project as open source at https://github.com/CLeARoboticsLab/MultiFidelityPolicyGradients.
@@
We build our implementation on top of the RL library Stable-Baselines3 (Raffin et al., 2021). For experimentation, we use random seeds {0, 1, 2} for hyperparameter tuning and {3, 4, 5, 6, 7} for evaluation. The High-Fidelity Only baseline is tuned using 3 seeds {0, 1, 2}. The resulting configuration serves as the shared backbone for the other methods, following standard practice in the off-dynamics RL literature (Xu et al., 2023; Lyu et al., 2024a). All additional hyperparameters are tuned with 2 seeds {0, 1}, either over more task settings (for the multi-fidelity methods) or longer training steps (for More High-Fidelity Data (15×)). All single-fidelity methods are tuned on the MuJoCo Hopper-v3, Walker2d-v3, and HalfCheetah-v3 tasks with original dynamics, chosen because they represent the midpoint among the varied dynamics settings. Multi-fidelity methods are tuned on the same three tasks under a representative subset of dynamics variations, namely $0.8\times$ and $2.0\times$ gravity. We verified that the tuned hyperparameters generalize well to other variation types and levels (cf. Appendix˜C). For the extra sensitive hyperparameters of PAR (reward-augmentation weight) and DARC (the standard deviation of Gaussian noise added to classifier inputs), we additionally tune them on the $0.8\times$ and $2.0\times$ friction variations of the three tasks. We note that this protocol grants the baselines more tuning privilege than MFPG, highlighting the simplicity and minimum tuning overhead of MFPG.
@@
High-Fidelity Only. We implement the standard REINFORCE algorithm (Williams, 1992) with state-value subtraction. Subject to the high-fidelity sample restrictions per policy update described in Section˜6.1, we perform grid search over the learning rate $\{1\times 10^{-4},2\times 10^{-4},5\times 10^{-4},7\times 10^{-4},9\times 10^{-4},1\times 10^{-3},2\times 10^{-3}\}$, discount factor $\gamma$ {0.95, 0.97, 0.98, 0.99, 0.995}, advantage normalization {True, False}, maximum gradient norm {0.5, 1, 2}, and state-value loss weight {0.25, 0.5, 1}. We adopt the default network architectures for the policy and state-value networks in Stable-Baselines3 (Raffin et al., 2021). The tuned REINFORCE configuration is then adopted as the shared algorithmic backbone for all multi-fidelity methods. The full hyperparameter configuration, along with those for all other methods, is provided in Table˜1.
@@
DARC. DARC (Eysenbach et al., ) trains two domain classifiers $q_{\theta_{\text{SAS}}}(\text{high}\mid s_{t},a_{t},s_{t+1})$ and $q_{\theta_{\text{SA}}}(\text{high}\mid s_{t},a_{t})$ to predict how likely an observed transition $(s_{t},a_{t},s_{t+1})$ or state-action pair $(s_{t},a_{t})$ comes from the high-fidelity environment. The classifiers are trained by minimizing cross-entropy losses over the replay buffers ${\mathcal{D}}_{\text{high}}$ and ${\mathcal{D}}_{\text{low}}$ of high- and low-fidelity samples, respectively:
@@
$\Delta r(s_{t},a_{t},s_{t+1})$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\log p(\text{high}\mid s_{t},a_{t},s_{t+1})}}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ {\log p(\text{high}\mid s_{t},a_{t})}}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\log p(\text{low}\mid s_{t},a_{t},s_{t+1})}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ {\log p(\text{low}\mid s_{t},a_{t})}}.$ (8)
@@
We adopt the standard implementation from the ODRL benchmark (https://github.com/OffDynamicsRL/off-dynamics-rl), cross-validated against the authors’ original code (https://github.com/google-research/google-research/tree/master/darc). To adapt DARC to the on-policy REINFORCE algorithm, we retain the original classifier-learning and reward-augmentation mechanisms, substituting only the policy optimization backbone. At each update, classifiers are trained using off-policy samples from the high- and low-fidelity replay buffers, as in the original implementation, by minimizing the cross-entropy losses in Eq.˜7. The learned classifiers are then used to augment the rewards, cf., Eq.˜8, for low-fidelity, on-policy samples, which are subsequently employed to compute the REINFORCE updates and learn the state-value function.
@@
PAR. Similar in spirit to DARC, PAR (Lyu et al., 2024a) also estimates the dynamics gap of low-fidelity transitions for reward augmentation. However, instead of domain classification, PAR adopts a representation-learning approach: it trains a state encoder $z_{1,t}=f_{\zeta}(s_{t})$ and a state–action encoder $z_{2,t}=g_{\nu}(z_{1,t},a_{t})$ by minimizing a latent dynamics consistency loss over the high-fidelity replay buffer:
@@
We adopt the standard implementation from the ODRL benchmark (https://github.com/OffDynamicsRL/off-dynamics-rl), cross-validated against the authors’ original code (https://github.com/dmksjfl/PAR). Analogous to our adaptation of DARC to the on-policy REINFORCE algorithm, we adapt PAR by retaining its original representation-learning and reward-augmentation mechanisms while substituting only the policy optimization backbone. At each update, we draw off-policy, high-fidelity samples from the replay buffer to train the encoders by minimizing Eq.˜9. The trained encoders are then applied to augment rewards for on-policy low-fidelity samples, as in Eq.˜10.
@@
Figure 14: Sweep of the reward-augmentation weight $\beta$ for PAR on two tasks. The relative performance across $\beta$ values varies substantially between tasks.
Unlike DARC, which relies solely on augmented low-fidelity samples for policy and value function learning, PAR leverages both on-policy high-fidelity and augmented low-fidelity samples for REINFORCE policy updates and state-value function learning, consistent with the original PAR algorithm.
@@
On top of the shared REINFORCE backbone, we further tune the reward-augmentation weight $\beta$ in Eq.˜10 over the grid {0.1, 0.5, 1, 2}, as PAR is particularly sensitive to this parameter; cf. Fig.˜14. The original PAR (Lyu et al., 2024a) employs a per-task weight for this reason, and we follow the same convention, reporting the selected $\beta$ for each task in Table˜1. We note, however, that this requirement introduces a computationally expensive adaptation overhead, unlike other algorithms, which do not require per-task tuning. For other hyperparameters—such as encoder architecture, learning rate, and batch size—we adopt the defaults from the original paper (Lyu et al., 2024a). The resulting encoder training curves are stable.
@@
MFPG. Building on the shared REINFORCE backbone, we perform a grid search over the exponential moving-average weight $\eta_{\text{ma}}$ in Eq.˜2 using values {0.92, 0.95, 0.99}. MFPG is found to be relatively robust to this hyperparameter, as illustrated in Section˜6.5. In off-dynamics task settings, we drop the control variate term $c(X^{\pi_{\theta}}_{\tau^{l}}-\hat{\mu}^{l})$ whenever the estimated correlation coefficient for the current batch, i.e., $\hat{\rho}(X^{\pi_{\theta}}_{\tau^{h}},X^{\pi_{\theta}}_{\tau^{l}})_{\text{new}}$ in Eq.˜2, is negative. Since high- and low-fidelity samples should in principle be positively correlated in off-dynamics settings, a negative correlation suggests either low-quality samples or noise in correlation estimation. In practice, such cases occur only rarely. This engineering trick helps mitigate noise in estimating the control variate coefficient, which can arise from extreme low-fidelity samples. Developing more principled estimation schemes remains an orthogonal direction for future work.
@@
Low-Fidelity Only (100×). The Low-Fidelity Only baseline serves as an ablation of high-fidelity samples from the multi-fidelity approaches. Accordingly, we adopt the same hyperparameter configuration as the shared REINFORCE backbone, except that each policy and state-value function update is performed using abundant low-fidelity data in place of limited high-fidelity samples.
@@
More High-Fidelity Data (15×). To ensure fair benefit from the additional high-fidelity samples, we tune the sensitive hyperparameters for the oracle baseline More High-Fidelity Data (15×) separately from the shared REINFORCE backbone. Specifically, we tune the learning rate over the grid $\{5\times 10^{-4},7\times 10^{-4},9\times 10^{-4},1\times 10^{-3},2\times 10^{-3},4\times 10^{-3}\}$ and the batch size (i.e., samples per policy update) over the grid {1024, 2048, 3072, 4096, 6144}. Because of the expensive, longer training time for this baseline, we first run all configurations for 3 million steps, then select the best 7 configurations and extend them to 15 million steps to determine the final configuration. Other hyperparameters are kept consistent with the shared REINFORCE backbone, as they are relatively insensitive.
@@
Low-fidelity data amount. In baseline comparisons, we assume that low-fidelity samples are cheap to generate, and all multi-fidelity methods may supplement limited high-fidelity data with up to $100\times$ additional low-fidelity samples per policy update. While PAR has been empirically observed to benefit from more low-fidelity data in some task settings (Lyu et al., 2024a), we also found it to be sensitive to the amount of low-fidelity data in certain cases. To maximize baseline performance, we additionally tune the ratio of low- to high-fidelity samples per policy update over the grid $\{10\times,20\times,50\times,100\times\}$ for PAR and DARC. For DARC, performance is generally maximized with $100\times$ low-fidelity samples. For PAR, performance is maximized across tasks at $20\times$. The Low-Fidelity Only baseline is assigned $100\times$ low-fidelity samples to match the budget. We implement parallelized low-fidelity environments, each generating $10\times$ low-fidelity samples per policy update; hence, DARC and Low-Fidelity Only employ 10 environments, while PAR uses 2 environments. For MFPG, we did not tune the amount of low-fidelity samples, since MFPG is in theory stable with respect to low-fidelity sample amounts, so long as the low-fidelity sample amount is sufficient to estimate the low-fidelity sample mean $\hat{\mu}^{l}$. Our empirical results in Fig.˜9 support this point. Therefore, instead of using 10 parallel low-fidelity environments (which would exceed the $100\times$ budget, given that MFPG also requires $1\times$ correlated low-fidelity samples), we adopt 9 parallel environments in all main results, corresponding to $90\times$ low-fidelity samples (plus $1\times$ correlated low-fidelity samples). In some auxiliary variance and sensitivity analysis in Sections˜6.2 and 6.5, we use $100\times$ (uncorrelated) low-fidelity samples; we specify the low-fidelity sample amount separately in those sections. Finally, we emphasize that DARC and PAR receive more tuning budget than MFPG, underscoring the simplicity and low adaptation overhead of MFPG.
@@
Complete hyperparameter configuration. Table˜1 reports the full hyperparameter configuration for all methods in the baseline comparison results. As noted above, the High-Fidelity Only configuration serves as the shared backbone for other methods, with algorithm-specific hyperparameters listed on top of this backbone. Blue entries indicate hyperparameters that replace the corresponding values in the backbone.
@@
Reward augmentation weight $\beta$ 0.5 (Hopper-v3), 0.1 (Walker2d-v3), 1.0 (HalfCheetah-v3)

### v4

@@ Appendix F Implementation Details and Hyperparameter Setup
This section provides implementation details and the hyperparameter setup for all evaluated approaches. Details of the off-dynamics task settings described in Section˜6.1 are available in the ODRL benchmark (Lyu et al., 2024b). Table˜1 summarizes the hyperparameter configurations for all evaluated methods. To ensure reproducibility, we include our code in the supplementary materials and will release the project as open source upon acceptance.
@@
We build our implementation on top of the RL library Stable-Baselines3 (Raffin et al., 2021). For experimentation, we use random seeds {0, 1, 2} for hyperparameter tuning and {3, 4, $\dots$ , 22} for evaluation in Sections˜6.3, 6.4 and 6.5 ({3, 4, $\dots$ , 7} for the additional 24 tasks in Appendix˜E). The High-Fidelity Only baseline is tuned using 3 seeds {0, 1, 2}. The resulting configuration serves as the shared backbone for the other methods, following standard practice in the off-dynamics RL literature (Xu et al., 2023; Lyu et al., 2024a). All additional hyperparameters are tuned with 2 seeds {0, 1}, either over more task settings (for the multi-fidelity methods) or longer training steps (for More High-Fidelity Data (15×)). All single-fidelity methods are tuned on the MuJoCo Hopper-v3, Walker2d-v3, and HalfCheetah-v3 tasks with original dynamics, chosen because they represent the midpoint among the varied dynamics settings. Multi-fidelity methods are tuned on the same three tasks under a representative subset of dynamics variations, namely $0.8\times$ and $2.0\times$ gravity. We verified that the tuned hyperparameters generalize well to other variation types and levels (cf. Appendix˜E). For the extra sensitive hyperparameters of PAR (reward-augmentation weight) and DARC (the standard deviation of Gaussian noise added to classifier inputs), we additionally tune them on the $0.8\times$ and $2.0\times$ friction variations of the three tasks. We note that this protocol grants the baselines more tuning privilege than MFPG, highlighting the simplicity and minimum tuning overhead of MFPG.
High-Fidelity Only.
@@
We implement the standard REINFORCE algorithm (Williams, 1992) with state-value subtraction. Subject to the high-fidelity sample restrictions per policy update described in Section˜6.1, we perform grid search over the learning rate $\{1\times 10^{-4},2\times 10^{-4},5\times 10^{-4},7\times 10^{-4},9\times 10^{-4},1\times 10^{-3},2\times 10^{-3}\}$, discount factor $\gamma$ {0.95, 0.97, 0.98, 0.99, 0.995}, advantage normalization {True, False}, maximum gradient norm {0.5, 1, 2}, and state-value loss weight {0.25, 0.5, 1}. We adopt the default network architectures for the policy and state-value networks in Stable-Baselines3 (Raffin et al., 2021). The tuned REINFORCE configuration is then adopted as the shared algorithmic backbone for all multi-fidelity methods. The full hyperparameter configuration, along with those for all other methods, is provided in Table˜1.
DARC.
@@
DARC (Eysenbach et al., ) trains two domain classifiers $q_{\theta_{\text{SAS}}}(\text{high}\mid s_{t},a_{t},s_{t+1})$ and $q_{\theta_{\text{SA}}}(\text{high}\mid s_{t},a_{t})$ to predict how likely an observed transition $(s_{t},a_{t},s_{t+1})$ or state-action pair $(s_{t},a_{t})$ comes from the high-fidelity environment. The classifiers are trained by minimizing cross-entropy losses over the replay buffers ${\mathcal{D}}_{\text{high}}$ and ${\mathcal{D}}_{\text{low}}$ of high- and low-fidelity samples, respectively:
@@
$\Delta r(s_{t},a_{t},s_{t+1})$ $={\color[rgb]{0,0,0}{\log p(\text{high}\mid s_{t},a_{t},s_{t+1})}}-{\color[rgb]{0,0,0}\ {\log p(\text{high}\mid s_{t},a_{t})}}-{\color[rgb]{0,0,0}{\log p(\text{low}\mid s_{t},a_{t},s_{t+1})}}+{\color[rgb]{0,0,0}\ {\log p(\text{low}\mid s_{t},a_{t})}}.$ (10)
@@
We adopt the standard implementation from the ODRL benchmark (https://github.com/OffDynamicsRL/off-dynamics-rl), cross-validated against the authors’ original code (https://github.com/google-research/google-research/tree/master/darc). To adapt DARC to the on-policy REINFORCE algorithm, we retain the original classifier-learning and reward-augmentation mechanisms, substituting only the policy optimization backbone. At each update, classifiers are trained using off-policy samples from the high- and low-fidelity replay buffers, as in the original implementation, by minimizing the cross-entropy losses in Eq.˜9. The learned classifiers are then used to augment the rewards, cf., Eq.˜10, for low-fidelity, on-policy samples, which are subsequently employed to compute the REINFORCE updates and learn the state-value function.
@@
PAR.
Similar in spirit to DARC, PAR (Lyu et al., 2024a) also estimates the dynamics gap of low-fidelity transitions for reward augmentation. However, instead of domain classification, PAR adopts a representation-learning approach: it trains a state encoder $z_{1,t}=f_{\zeta}(s_{t})$ and a state–action encoder $z_{2,t}=g_{\nu}(z_{1,t},a_{t})$ by minimizing a latent dynamics consistency loss over the high-fidelity replay buffer:
@@
Figure 18: Sweep of the reward-augmentation weight $\beta$ for PAR on two tasks. The relative performance across $\beta$ values varies substantially between tasks. Bar heights indicate medians, and error bars denote the minimum and maximum values.
We adopt the standard implementation from the ODRL benchmark (https://github.com/OffDynamicsRL/off-dynamics-rl), cross-validated against the authors’ original code (https://github.com/dmksjfl/PAR). Analogous to our adaptation of DARC to the on-policy REINFORCE algorithm, we adapt PAR by retaining its original representation-learning and reward-augmentation mechanisms while substituting only the policy optimization backbone. At each update, we draw off-policy, high-fidelity samples from the replay buffer to train the encoders by minimizing Eq.˜11. The trained encoders are then applied to augment rewards for on-policy low-fidelity samples, as in Eq.˜12. Unlike DARC, which relies solely on augmented low-fidelity samples for policy and value function learning, PAR leverages both on-policy high-fidelity and augmented low-fidelity samples for REINFORCE policy updates and state-value function learning, consistent with the original PAR algorithm.
@@
On top of the shared REINFORCE backbone, we further tune the reward-augmentation weight $\beta$ in Eq.˜12 over the grid {0.1, 0.5, 1, 2}, as PAR is particularly sensitive to this parameter; cf. Fig.˜18. The original PAR (Lyu et al., 2024a) employs a per-task weight for this reason, and we follow the same convention, reporting the selected $\beta$ for each task in Table˜1. We note, however, that this requirement introduces a computationally expensive adaptation overhead, unlike other algorithms, which do not require per-task tuning. For other hyperparameters—such as encoder architecture, learning rate, and batch size—we adopt the defaults from the original paper (Lyu et al., 2024a). The resulting encoder training curves are stable.
MFPG.
@@
Building on the shared REINFORCE backbone, we perform a grid search over the exponential moving-average weight $\eta_{\text{ma}}$ in Eq.˜3 using values {0.92, 0.95, 0.99}. We find that MFPG ’s median performance consistently exceeds that of the High-Fidelity Only baseline across the tested values of $\eta_{\text{ma}}$, even though the absolute performance metrics vary with $\eta_{\text{ma}}$ . In off-dynamics task settings, we drop the control variate term $c(X^{\pi_{\theta}}_{\tau^{l}}-\hat{\mu}^{l})$ whenever the estimated correlation coefficient for the current batch, i.e., $\hat{\rho}^{\text{batch}}_{k}(X_{\tau^{h}}^{\pi_{\theta_{k}}},X_{\tau^{l}}^{\pi_{\theta_{k}}})$ in Eq.˜3, is negative. Since high- and low-fidelity samples should in principle be positively correlated in off-dynamics settings, a negative correlation suggests either low-quality samples or noise in correlation estimation. In practice, such cases occur only rarely. This engineering trick helps mitigate noise in estimating the control variate coefficient, which can arise from extreme low-fidelity samples. Developing more principled estimation schemes remains an orthogonal direction for future work.
Low-Fidelity Only (100×).
@@
The Low-Fidelity Only baseline serves as an ablation of high-fidelity samples from the multi-fidelity approaches. Accordingly, we adopt the same hyperparameter configuration as the shared REINFORCE backbone, except that each policy and state-value function update is performed using abundant low-fidelity data in place of limited high-fidelity samples.
More High-Fidelity Data (15×).
@@
To ensure fair benefit from the additional high-fidelity samples, we tune the sensitive hyperparameters for the oracle baseline More High-Fidelity Data (15×) separately from the shared REINFORCE backbone. Specifically, we tune the learning rate over the grid $\{5\times 10^{-4},7\times 10^{-4},9\times 10^{-4},1\times 10^{-3},2\times 10^{-3},4\times 10^{-3}\}$ and the batch size (i.e., samples per policy update) over the grid {1024, 2048, 3072, 4096, 6144}. Because of the expensive, longer training time for this baseline, we first run all configurations for 3 million steps, then select the best 7 configurations and extend them to 15 million steps to determine the final configuration. Other hyperparameters are kept consistent with the shared REINFORCE backbone, as they are relatively insensitive.
Low-fidelity data amount.
@@
In baseline comparisons, we assume that low-fidelity samples are cheap to generate, and all multi-fidelity methods may supplement limited high-fidelity data with up to $100\times$ additional low-fidelity samples per policy update. While PAR has been empirically observed to benefit from more low-fidelity data in some task settings (Lyu et al., 2024a), we also found it to be sensitive to the amount of low-fidelity data in certain cases. To maximize baseline performance, we additionally tune the ratio of low- to high-fidelity samples per policy update over the grid $\{10\times,20\times,50\times,100\times\}$ for PAR and DARC. For DARC, performance is generally maximized with $100\times$ low-fidelity samples. For PAR, performance is maximized across tasks at $20\times$. The Low-Fidelity Only baseline is assigned $100\times$ low-fidelity samples to match the budget. We implement parallelized low-fidelity environments, each generating $10\times$ low-fidelity samples per policy update; hence, DARC and Low-Fidelity Only employ 10 environments, while PAR uses 2 environments. For MFPG, we did not tune the amount of low-fidelity samples, since MFPG is in theory stable with respect to low-fidelity sample amounts, so long as the low-fidelity sample amount is sufficient to estimate the low-fidelity sample mean $\hat{\mu}^{l}$. Our empirical results in Fig.˜9 support this point. Therefore, instead of using 10 parallel low-fidelity environments (which would exceed the $100\times$ budget, given that MFPG also requires $1\times$ correlated low-fidelity samples), we adopt 9 parallel environments in all main results, corresponding to $90\times$ low-fidelity samples (plus $1\times$ correlated low-fidelity samples). In some auxiliary variance and sensitivity analysis in Sections˜6.2 and 6.5, we use $100\times$ (uncorrelated) low-fidelity samples; we specify the low-fidelity sample amount separately in those sections. Finally, we emphasize that DARC and PAR receive more tuning budget than MFPG, underscoring the simplicity and low adaptation overhead of MFPG.
Complete hyperparameter configuration.
@@
Table˜1 reports the full hyperparameter configuration for all methods in the baseline comparison results. As noted above, the High-Fidelity Only configuration serves as the shared backbone for other methods, with algorithm-specific hyperparameters listed on top of this backbone. Brown entries indicate hyperparameters that replace the corresponding values in the backbone.
@@
Reward augmentation weight $\beta$ 0.5 (Hopper), 0.1 (Walker2d), 1.0 (HalfCheetah)
@@
Appendix G Extending MFPG to Broader Algorithms and Settings
For the first future-work direction discussed in Section˜7, we highlight two subdirections. First, a natural next step is to extend the proposed multi-fidelity control variate approach to a broader class of modern RL algorithms.
A concrete starting point in this direction is modern on-policy actor-critic methods, such as proximal policy optimization (PPO; Schulman et al., 2017). As discussed in Section˜4, the MFPG framework can, in principle, be instantiated with other on-policy policy gradient methods by changing how the random variable $X_{\tau}^{\pi_{\theta}}$ is computed from sampled trajectories. In this work, we focus on the classic REINFORCE algorithm to simplify the theoretical analysis and establish convergence guarantees. Modern actor-critic methods, however, deliberately trade bias for lower variance in their policy gradient estimates, e.g. via generalized advantage estimation (Schulman et al., 2015). As a result, the high-fidelity policy gradient estimator is typically biased (violating Assumption˜5 in Section˜5), and extending our convergence analysis to this setting is significantly more challenging. Nevertheless, studying the bias–variance trade-off in this regime may still allow MFPG to yield meaningful variance reduction and performance gain for contemporary on-policy actor-critic methods, and thus merits deeper investigation. Moreover, within the on-policy setting, the proposed MFPG framework could, in principle, be extended in a model-based fashion, where the low-fidelity control-variate random variable $X_{\tau^{l}}^{\pi_{\theta}}$ and its sample mean $\hat{\mu}^{l}$ are computed from a learned dynamics model.
Another important direction is to extend MFPG to off-policy settings. On the one hand, the MFPG framework might be adapted to importance-sampling-based off-policy policy-gradient estimators, e.g. by replacing the high- and low-fidelity random variables with off-policy, importance-weighted policy-gradient estimators. In such settings, variance reduction is arguably even more critical than in on-policy policy-gradient methods (Metelli et al., 2020), making multi-fidelity control variates particularly appealing. On the other hand, if one wishes to extend an MFPG-type approach to modern off-policy algorithms such as soft actor-critic (SAC) (Haarnoja et al., 2018), developing multi-fidelity control variate techniques for off-policy temporal-difference (TD) learning would be an important step. This extension is less direct. In contrast to on-policy policy-gradient methods, where high-variance Monte Carlo returns motivate sophisticated variance-reduction techniques, most modern off-policy algorithms, such as SAC, rely on one-step TD learning and experience replay, which tend to produce relatively low-variance value targets. Their main challenges are primarily bias-related, e.g. value overestimation (Fujimoto et al., 2018), distribution shift (Kumar et al., 2019), and bias accumulation over long horizons (Park et al., 2025). Even so, multi-fidelity control variates may still help reduce noise in bootstrapped targets and the variance of gradient estimates, e.g. in the policy-improvement (policy extraction) steps. In particular, pathwise gradient estimators used in TD-learning–based approaches can still suffer from high variance, depending on the conditioning and smoothness of the learned value function (Mohamed et al., 2020).
A separate line of future work is to extend the multi-fidelity control variate technique to incorporate offline datasets. One scenario is that a large volume of low-fidelity data is available as a static offline dataset, while a smaller amount of high-fidelity data can be gathered online. In principle, one could develop a mechanism similar to the MFPG framework in this work for incorporating such offline (and therefore off-policy) low-fidelity data in a way that aims not to introduce additional bias into the policy gradients estimated from online high-fidelity samples. Concretely, the low-fidelity sample mean $\hat{\mu}^{l}$ would be estimated from an offline dataset. If, in addition, the offline dataset logs (or allows reconstruction of) the noise samples used for action selection, then one could draw correlated high-fidelity trajectories in the high-fidelity environment via the reparameterization scheme introduced in Section˜4 and use the offline low-fidelity trajectories to construct control variates. Another scenario is the standard offline RL setting, in which the high-fidelity samples themselves come from an offline dataset that may be augmented with online low-fidelity simulation data. Extending MFPG to this setting is less direct. The challenges are at least two-fold. First, the specific MFPG framework studied in this work requires the high-fidelity random variable $X_{\tau^{h}}^{\pi_{\theta}}$ to be computed from on-policy samples, whereas offline data are inherently off-policy. One might use importance sampling to reweight the offline data and then use online low-fidelity simulation data to reduce the variance of the resulting importance-weighted high-fidelity gradient estimates. Second, extending multi-fidelity control variates to commonly used, TD-learning-based offline RL algorithms (Kumar et al., 2020; Kostrikov et al., 2021) faces challenges similar to those discussed above for off-policy TD methods: the dominant issues are bias-related, e.g. inaccurate value estimates for out-of-distribution actions and severe distribution shift between the behavior and target policies (Levine et al., 2020).